OpenStudy (anonymous):

Is this correct? Derivative of y = sqrt[(x^2+1)(sin3(x)]

OpenStudy (anonymous):

y' = (x^2+1)^1/2(3/2)(sinx)(cosx)+(sinx)^3/2 (1/2)(x^2+1)^-1/2 (2x)

OpenStudy (anonymous):

im confused haha sorry

zepdrix (zepdrix):

Hmm I can't seem to figure out what you did bobo :( it's kind of hard to read.

OpenStudy (anonymous):

Can you teach me the right way?

zepdrix (zepdrix):

\[\Large y=\left[(x^2+1)\sin3x\right]^{1/2}\] Applying Power Rule first,\[\Large y'=\frac{1}{2}\left[(x^2+1)\sin3x\right]^{-1/2}\color{royalblue}{\left[(x^2+1)\sin3x\right]'}\] Chain Rule tells us that we need to take a copy of the inner function and multiply by its derivative. So we still need to take the derivative of this blue portion. Understand the power rule step?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Then looks like we need to apply the product rule to the blue part, giving us this setup:\[\large y'=\frac{1}{2}\left[(x^2+1)\sin3x\right]^{-1/2}\left[\color{#CC0033}{(x^2+1)'}\sin3x+(x^2+1)\color{#CC0033}{(\sin3x)'}\right]\]

OpenStudy (anonymous):

Okay I follow that

zepdrix (zepdrix):

From here, you need to take the derivative of the red parts. What do you get for those? :x

OpenStudy (anonymous):

2x and cos3x?

zepdrix (zepdrix):

Hmm close! Don't forget the chain rule on your sin3x. It should give us an extra 3, right?\[\large y'=\frac{1}{2}\left[(x^2+1)\sin3x\right]^{-1/2}\left[\color{royalblue}{(2x)}\sin3x+(x^2+1)\color{royalblue}{(3\cos3x)}\right]\]

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

is there not something that comes after?

zepdrix (zepdrix):

Mmmm no nothing else \c:/

OpenStudy (anonymous):

Thank you very much @zepdrix You're a great teacher

zepdrix (zepdrix):

yay team :D